A rogue physics teacher with a mass of 53 kg is rolling on a track with a speed of 3 m/s. If friction does -100 J of work on the teacher, how much impulse did it impart?
Answer #1

here,

the mass of Teacher , m = 53 kg

initial speed , u = 3 m/s

let the final speed be v

using Work-energy theorm

0.5 *m * ( v^2 - u^2) = Work done by friction force

0.5 * 53 * ( v^2 - 3^2) = - 100

v = 2.29 m/s

the impulse delevired , I = m * ( v - u)

I = 53 * ( 2.29 - 3) kg.m/s

I = - 37.8 kg.m/s